OpenStudy (anonymous):

how do you solve this problem 3(x-4)=-x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

3x-12=-x 4x=12 x=3

OpenStudy (anonymous):

distribute the ex. then subtract x from left. then factor out the x.

OpenStudy (anonymous):

you do distributive property

OpenStudy (anonymous):

3x-12= -x 3x+x= 12 4x= 12 x= 3

OpenStudy (anonymous):

3x-12=-x subtract 3x from -12 -12=-x-3x -12=x(-1-3) then divide. (-1-3) x=-12/(-4) x=3

OpenStudy (anonymous):

(-1-3)=-4